# Combining Constrained Diffusion Models and Numerical Solvers for Efficient and Robust Non-Convex Trajectory Optimization

### v3

@@ Abstract
Motivated by the need to solve open-loop optimal control problems with computational efficiency and reliable constraint satisfaction, we introduce a general framework that combines diffusion models and numerical optimization solvers. Optimal control problems are rarely solvable in closed form, hence they are often transcribed into numerical trajectory optimization problems, which then require initial guesses. These initial guesses are supplied in our framework by diffusion models. To mitigate the effect of samples that violate the problem constraints, we develop a novel constrained diffusion model to approximate the true distribution of locally optimal solutions with an additional constraint violation loss in training. To further enhance the robustness, the diffusion samples as initial guesses are fed to the numerical solver to refine and derive final optimal (and hence feasible) solutions. Experimental evaluations on three tasks verify the improved constraint satisfaction and computational efficiency with 4$\times$ to 30$\times$ acceleration using our proposed framework, which generalizes across trajectory optimization problems and scales well with problem complexity.
@@ 1 Introduction
Optimal control problems are fundamental in decision-making for autonomous agents including self-driving cars, robot manipulators, and spacecraft trajectories to name just a few. In the open-loop case, with a predetermined initial configuration, the goal is to plan an optimal path and control history for an agent to reach a target while satisfying the dynamics and safety constraints. These problems often have non-convex structure, which may be due to complex nonlinear system dynamics and environmental constraints. Ultimately, this results in multiple minima and solutions that may have different qualitative behavior with various tradeoffs. Identifying these diverse and locally optimal (and hence feasible) solutions efficiently remains a significant challenge in the field.
@@
The open-loop optimal control problem can be cast as a numerical trajectory optimization problem using a variety of approaches [4]. When the optimal control problem involves nonlinear functions, this often results in a nonlinear program (NLP) where an initial guess is required to get the solution. In the case where one is interested in solving for several solutions, a global search is needed. This is essentially a two-step process: (i). sampling a distribution on the control space, and (ii). using this sample as an initial guess to the NLP solver that hopefully generates a solution. For each solution from the NLP, optimality (and hence feasibility) is often verified by the solver by checking whether the first order necessary conditions, the Karush-Kuhn-Tucker (KKT) conditions, are satisfied [29, 25]. However, the high dimensionality and non-convex nature of the problem often make the two-step global search process time-consuming and cannot satisfy the efficiency demands of autonomous systems.
@@
Figure 1: Three trajectory optimization tasks in our experiments (from left to right): tabletop manipulation (161 DoF); two-car reach-avoid problem (81 DoF); cislunar transfer problem (64 DoF). The degree-of-freedom (DoF) indicates the dimension of control variables for each task.
@@
Machine learning has also gained popularity in trajectory optimization, like warm starting the convexified trajectory optimization problems such as Quadratic Programming (QP) [11, 9, 38]. These approaches primarily focus on predicting a single solution to the QP problem to improve the computational efficiency, rather than modeling the distribution of diverse solutions to the original non-convex problem. Recently, generative models, particularly diffusion models, have shown good performance in generating trajectories or controls from a distribution [24, 1, 12, 8, 42]. When conditioned on the environmental parameters, diffusion models can efficiently generate appropriate and diverse trajectories even in a priori unknown environment. However, the diffusion model output trajectories are not robust to prediction error and may violate the dynamic and safety constraints, which can cause catastrophic damage to safety-critical systems. It’s also difficult for the diffusion model to directly output locally optimal trajectories. While efforts have been made to integrate safety constraints into generative models [32, 10, 49, 48, 36, 6], it is still challenging to achieve solution quality, computational efficiency, diversity, robustness and feasibility.
@@
In this paper, we propose a general framework for efficient non-convex trajectory optimization with enhanced robustness for constraint satisfaction. This framework creates a synergy between a novel constrained diffusion model and a numerical optimization solver. We first formulate the trajectory optimization problem as a nonlinear program, where a numerical solver can be used to collect solutions as training data. Then we train a conditional diffusion model to efficiently sample a set of diverse and high-quality trajectory predictions, from which the solver can fine-tune to obtain final solutions with enhanced feasibility and optimality. We also developed an innovative constrained diffusion model that incorporates an additional constraint violation loss in the training process to minimize the violation in the samples.
@@
Our framework has the following appealing properties: (a). Computational Efficiency. Trained with offline solution data, the diffusion model can predict high-quality solutions that are close to local optima and warm-start the trajectory optimization. This significantly accelerates the solver convergence with considerable computational time savings. (b). Robustness (Constraint Satisfaction). We used a constrained diffusion model specifically designed to produce samples with minimal constraint violations. These sampled trajectories are then further refined by a numerical solver to enhance the feasibility and optimality. (c). Automatic Data Generation. With a numerical solver, we can automatically generate diverse problems and collect corresponding solutions, which allows for on-demand data augmentation whenever required. (d). Generalization. Our framework is suitable for trajectory optimization across various applications with flexible environment settings.
We demonstrate the efficacy of our framework on three non-convex trajectory optimization problems for open-loop optimal control: tabletop manipulation, two-car reach-avoid, and cislunar low thrust transfer for a spacecraft. Our framework improves the efficiency of obtaining solutions compared to the baseline while enhancing constraint satisfaction.
@@ 2 Related Work
Machine learning methods have been widely used to solve a variety of optimization problems with different techniques, including convex Quadratic Programming (QP) [50, 51, 11, 38], mixed-integer programming [9], combinatorial optimization [43, 23] and black-box optimization [30, 45, 28], etc. To improve the constraint satisfaction in optimization problems, neural networks have been incorporated with fully differentiable constraint function [16]. Conditional variational autoencoder (CVAE) and long short-term memory (LSTM) are proposed to warm-start the general nonlinear trajectory optimization problems [32]. Another branch of work leverages machine learning methods and pre-collected solution data points for solving partial differential equations (PDEs) [7]. One of the most pupular work is physics-informed neural networks (PINN) [37], which introduces the solution prediction model with a hybrid training loss for labeled solution prediction and violation of the original PDEs.
@@
The diffusion models [39, 41] is able to sample the multi-modal distribution with a stochastic differential equation (SDE). The models are further developed into denoising diffusion probabilistic models (DDPM) [21], denoising diffusion implicit models (DDIM) [40] for application of image generation. Both classifier-guided diffusion [13] and classifier-free guidance [22] are proposed to achieve conditional generation with diffusion models. Given the strong expressiveness of the diffusion model, its application has been extended across domains from image generation to reinforcement learning [46, 14, 15], motion generation [44], robotics [12], etc.
Diffusion-based Trajectory Generation.
Diffusion models are becoming popular in robotics for generating controls or trajectories [1, 24, 33, 42, 12, 34, 15]. Many efforts have been made to improve the diffusion model for creating safe trajectories for safety-critical systems. Control barrier functions [48, 6] and collision-avoidance kernels [10] are employed to guide diffusion models to sample feasible trajectory solutions. To adapt to new constraints in testing, prior work investigated composing diffusion models to improve the generalizability [36, 49]. However, solely relying on diffusion-sampled trajectories will inevitably face the robustness issue of constraint violations and hardly achieve locally optimal solutions. There has been an effort to combine diffusion models with an inference-based trajectory optimization framework [36, 35], but it is not parallelizable and requires extensive computational effort for complex problems. Our work guides the diffusion model toward low constraint violation in the output through a novel loss function and combines it with a fully parallelizable numerical optimization method that scales well.
@@ 3 Preliminaries
We consider a general open-loop optimal control problem: assume a known agent’s dynamical model, the predefined initial configuration, and the obstacle settings, we aim to plan an optimal trajectory and controls for the agent to reach a target while satisfying the dynamics and safety constraints. With the direct transcription [4], this trajectory optimization problem can be discretized as the following nonlinear program $\mathcal{P}_{y}$:
@@
When the problem parameter $y$ is fixed, $\mathcal{P}_{y}$ in Eq. (1) typically presents a non-convex optimization problem with multiple local optima. To solve this problem $\mathcal{P}_{y}$, a gradient-based numerical solver $\pi$ can be used. A local optimum $x^{*}$ to $\mathcal{P}_{y}$ is verified by the solver $\pi$ that satisfies the KKT conditions (i.e., no constraint violation). To find different $x^{*}$, the user first provides multiple initial guesses $x^{0}$ for the decision variables $x$ to the solver $\pi$, typically drawn from a uniform distribution or some heuristic approach. Then the solver $\pi$ performs an iterative, gradient-based optimization until a local optimum $x^{*}$ is found. The popular methods to conduct such gradient-based optimization in non-convex problems include the Interior Point Method (IPM) [47] and the Sequential Quadratic Programming (SQP) [5].
@@
As expressive generative models, the diffusion probabilistic models [39, 21, 41] generate samples from random noises through an iterative denoising process. The forward process (i.e., diffusion process) iteratively adds Gaussian noise to sample $x_{k}$ at the $k$-th step according to the following equation:
@@
where $\widebar{\alpha}_{k}=\prod_{k^{\prime}=1}^{K}(1-\beta_{k^{\prime}})$ and $\varepsilon\sim\mathcal{N}(0,\mathbf{I})$ is a random sample from a standard Gaussian. The data generation process is the reverse denoising process that transforms random noise to data sample:
@@
with initial noisy data sampled from a random Gaussian distribution $x_{K}\sim\mathcal{N}(0,\mathbf{I})$. $p_{\theta}$ is the parameterized sampling distribution, and it can be optimized through an alternative $\varepsilon_{\theta}(x_{k},k)$ which is to predict $\varepsilon$ injected for $x_{k}$, for every diffusion step $k$. For conditional generation modeling, a conditional variable $y$ is added to both processes $q(x_{k+1}|x_{k},y)$ and $p_{\theta}(x_{k-1}|x_{k},y)$, respectively. The classifier-free guidance [22] can be applied to further promote the conditional information, which learns both conditioned and unconditioned noise predictors as $\varepsilon_{\theta}(x_{k},k,y)$ and $\varepsilon_{\theta}(x_{k},k,\varnothing)$. Specifically, $\varepsilon_{\theta}$ is optimized with the following loss function:
@@ 4 Methodology
In this section, we first proposed the general framework for efficient and robust non-convex trajectory optimization using diffusion models and numerical solvers. Then we introduce a novel constrained diffusion model specifically designed to guide the diffusion models to generate feasible solutions.
4.1 A General Trajectory Optimization Framework
Figure 2: The proposed general framework for non-convex trajectory optimization with a constrained diffusion model and a numerical solver.
@@ 4.1 A General Trajectory Optimization Framework
It has been observed that for similar trajectory optimization problems $\mathcal{P}_{y}$ in Eq. (1), their solutions $x^{*}$ often exhibit similar structures [2, 32]. Similar problems are those $\mathcal{P}_{y}$ with the same objective $J$ and constraints $g,h$, but with different parameters $y$. Thus, the key insight of our framework is to use a diffusion model to learn $p(x^{*}|y)$, the conditional probability distribution of locally optimal solutions, based on pre-solved similar problems. Then the diffusion model can generalize and predict diverse solutions $\tilde{x}_{\text{new}}^{*}$ to new scenarios where the condition $y_{\text{new}}$ value is unseen.
However, diffusion models inevitably make prediction errors, which can result in constraint violations. To address this, we use diffusion samples $\tilde{x}_{\text{new}}^{*}$ as initial guesses for the numerical solver, which fine-tunes them to derive the final solutions $x_{\text{new}}^{*}$, increasing both feasibility and optimality. This approach acts like a safety net, enhancing the robustness of the solutions.
@@
We illustrate our proposed data-driven trajectory optimization framework with diffusion models in Fig. 2. In the offline data generation process, we use a numerical solver $\pi$ to collect locally optimal solutions $x^{*}$ for various problem instances $\mathcal{P}_{y}$ with uniformly sampled initial guesses. The solutions with sub-optimal objective values are filtered out manually and the remaining dataset is used as the training data. We train a conditional diffusion model to learn $p(x^{*}|y)$ with classifier-free guidance [22]. In the online testing process when a new problem $\mathcal{P}_{y_{\text{new}}}$ is presented with a priori unknown value $y_{\text{new}}$, we first use the diffusion model to predict diverse solution candidates $\tilde{x}_{\text{new}}^{*}$ as initial guesses. Then the numerical solver $\pi$ only requires minor adjustments to derive the final solutions $x_{\text{new}}^{*}$ with improved feasibility and optimality. This warm-starting approach is fully parallelizable for both diffusion sampling and the solving process of $\pi$.
4.2 Constrained Diffusion Model

### v4

# DiffuSolve: Diffusion-based Solver for Non-convex Trajectory Optimization

@@ Abstract
Optimal trajectory design is computationally expensive for nonlinear and high-dimensional dynamical systems. The challenge arises from the non-convex nature of the optimization problem with multiple local optima, which usually requires a global search. Traditional numerical solvers struggle to find diverse solutions efficiently without appropriate initial guesses. In this paper, we introduce DiffuSolve, a general diffusion model-based solver for non-convex trajectory optimization. An expressive diffusion model is trained on pre-collected locally optimal solutions and efficiently samples initial guesses, which then warm-starts numerical solvers to fine-tune the feasibility and optimality. We also present DiffuSolve+, a novel constrained diffusion model with an additional loss in training that further reduces the problem constraint violations of diffusion samples. Experimental evaluations on three tasks verify the improved robustness, diversity, and a 2$\times$ to 11$\times$ increase in computational efficiency with our proposed method, which generalizes well to trajectory optimization problems of varying challenges.
@@ 1 Introduction
Optimal trajectory design is fundamental in decision-making for autonomous agents. In the open-loop case, with a predetermined initial configuration, the goal is to plan an optimal path and control sequence for an agent to reach a target while satisfying the dynamics and safety constraints. In the case of complex nonlinear system dynamics and environmental constraints, these problems often have non-convex structures. For example, within a chaotic dynamical system like the Circular Restricted Three-Body Problem (CR3BP) (Koon et al., 2000), the trajectory optimization problem for the spacecraft has multiple local optima of different qualitative behaviors with various tradeoffs (Hartmann et al., 1998; Russell, 2007). Identifying these diverse, locally optimal, and feasible solutions efficiently remains a significant challenge in the field.
@@
Control transcription of the open-loop trajectory optimization problem involving nonlinear functions results in a nonlinear program (NLP), where an initial guess is required to get the solution (Betts, 1998). When one is interested in solving for several solutions, a two-step global search is needed: (i) sampling a distribution on the control space, and (ii) using this sample as an initial guess to the NLP solver that hopefully generates a solution. For each solution from the NLP, optimality (and also feasibility) is often verified by the solver by checking whether the first-order necessary conditions, the Karush-Kuhn-Tucker (KKT) conditions, are satisfied (Kuhn & Tucker, 2013; Karush, 1939). However, the high dimensionality and non-convex nature of the problem often make the two-step global search process time-consuming such that the efficiency demands of autonomous systems are not met.
@@
Figure 1: Three trajectory optimization tasks in our experiments (from left to right): two-car reach-avoid problem; quadrotor navigation problem; cislunar transfer problem. The dimensions of the decision variables are 81, 241, and 64, respectively.
@@
Machine learning has also gained popularity in trajectory optimization, but primarily focuses on solving the convexified problem rather than learning the solution distribution to the original non-convex problem. For example, investigations have been made on warm-starting Quadratic Programs (QP) using neural networks (Chen et al., 2022; Cauligi et al., 2021; Sambharya et al., 2023). Recently, generative models, particularly diffusion models, have shown good performance in generating trajectories or controls from a distribution (Janner et al., 2022; Ajay et al., 2022; Chi et al., 2023; Carvalho et al., 2023; Sridhar et al., 2023). When conditioned on the environmental parameters, diffusion models can efficiently generate appropriate and diverse trajectories even in an a priori unknown environment. However, the diffusion model alone without the NLP solver may output trajectories that violate the dynamic and safety constraints, causing catastrophic damage to safety-critical systems. It also lacks the ability to verify the optimality of the sampled trajectories. While efforts have been made to integrate safety constraints into generative models (Li et al., 2023; Chang et al., 2023; Yang et al., 2023; Xiao et al., 2023; Power et al., 2023; Botteghi et al., 2023), it is still challenging to achieve solution quality, diversity, feasibility, and computational efficiency at the same time.
@@
In this paper, we propose DiffuSolve: a diffusion model-based solver for non-convex trajectory optimization with improved efficiency, diversity, and robustness. Our contribution is two-fold: First, we train a conditional diffusion model to learn the complex solution distribution of non-convex optimization problems, and efficiently sample a set of diverse and high-quality trajectory predictions. These samples are then used as warm starts for an NLP solver to derive locally optimal and feasible solutions, ensuring the robustness of DiffuSolve. Second, to mitigate the effect of diffusion samples that violate the problem constraints, we developed DiffuSolve+ with a novel constrained diffusion model that incorporates an additional constraint violation loss in the training process.
@@
DiffuSolve has the following appealing properties: (a). Computational Efficiency. The diffusion model samples high-quality solutions that are close to local optima as warm starts, from which the NLP solver can quickly converge. (b). Robustness and Diversity. A constrained diffusion model produces diverse samples with minimal constraint violations, further refined by an NLP solver for enhanced feasibility and optimality. (c). Automatic Data Generation. With an NLP solver, we can generate diverse problems and collect corresponding solutions ourselves, allowing for on-demand data augmentation whenever required. (d). Generalization. Our framework is suitable for trajectory optimization across various system dynamics and environment settings, with the potential for optimization problems in other areas like finance (Sambharya et al., 2023), etc.
We demonstrate the efficacy of our method on three non-convex trajectory optimization problems: two-car reach-avoid, quadrotor navigation, and cislunar low-thrust transfer for a spacecraft.
@@ 2 Related Work
Machine learning methods have been widely used to solve a variety of optimization problems with different techniques, including convex Quadratic Programming (QP) (Zeilinger et al., 2011; Zhang et al., 2019; Chen et al., 2022; Sambharya et al., 2023), mixed-integer programming (Cauligi et al., 2021), combinatorial optimization (Sun & Yang, 2023; Huang et al., 2023) and black-box optimization (Kumar & Levine, 2020; Trabucco et al., 2022; Krishnamoorthy et al., 2023), etc. For the Combinatorial Optimization problem, graph-based diffusion models are used to directly generate solutions (Sun & Yang, 2023; Huang et al., 2023). To improve the constraint satisfaction in optimization problems, neural networks have been incorporated with fully differentiable constraint function (Donti et al., 2021). Conditional variational autoencoder (CVAE) and long short-term memory (LSTM) are proposed to warm-start the general nonlinear trajectory optimization problems (Li et al., 2023). Another branch of work leverages machine learning methods and pre-collected solution data points for solving partial differential equations (PDEs) (Cai et al., 2021). One of the most pupular work is physics-informed neural networks (PINN) (Raissi et al., 2019), which introduces the solution prediction model with a hybrid training loss for labeled solution prediction and violation of the original PDEs.
@@
The diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2020b) is able to sample the multi-modal distribution with a stochastic differential equation (SDE). The models are further developed into denoising diffusion probabilistic models (DDPM) (Ho et al., 2020), denoising diffusion implicit models (DDIM) (Song et al., 2020a), etc., for application of image generation. Both classifier-guided diffusion (Dhariwal & Nichol, 2021) and classifier-free guidance (Ho & Salimans, 2022) are proposed to achieve conditional generation with diffusion models.
Diffusion-based Trajectory Generation with Constraints.
Diffusion models are becoming popular in robotics for generating controls or trajectories (Ajay et al., 2022; Janner et al., 2022; Liang et al., 2023; Sridhar et al., 2023; Chi et al., 2023; Mishra et al., 2023; Ding et al., 2024). Many efforts have been made to improve the diffusion model for creating safe trajectories for safety-critical systems. Control barrier functions (Xiao et al., 2023; Botteghi et al., 2023) and collision-avoidance kernels (Chang et al., 2023) are employed to guide diffusion models to sample feasible trajectory solutions. To adapt to new constraints in testing, prior work investigated composing diffusion models to improve the generalizability (Power et al., 2023; Yang et al., 2023).
Different from existing work, our paper focuses on high-dimensional, highly nonlinear trajectory optimization problems with many local optima, and aims to find diverse solutions efficiently. Instead of incorporating constraints in the sampling process, our constrained diffusion model in DiffuSolve+ includes a novel loss in training, which can be combined with the current work to further reduce the violation. Moreover, DiffuSolve and DiffuSolve+ are equipped with an NLP solver that finetunes the diffusion samples until feasibility and optimality are obtained at its best.
@@ 3 Preliminaries
We consider a general open-loop trajectory optimization problem: assume a known agent’s dynamical model, the predefined initial configuration, and the obstacle settings, we aim to plan an optimal trajectory and controls for the agent to reach a target while satisfying the dynamics and safety constraints. With the direct transcription (Betts, 1998), this trajectory optimization problem can be discretized as the following nonlinear program $\mathcal{P}_{y}$:
@@
When the problem parameter $y$ is fixed, $\mathcal{P}_{y}$ in Eq. (1) typically presents a non-convex optimization problem with multiple local optima. To solve this problem $\mathcal{P}_{y}$, a gradient-based NLP solver $\pi$ can be used. A local optimum (feasible) $x^{*}$ to $\mathcal{P}_{y}$ is verified by the solver $\pi$ that satisfies the KKT conditions with no constrained violation. To find different $x^{*}$, the user first provides multiple initial guesses $x^{0}$ for the decision variables $x$ to the solver $\pi$, typically drawn from a uniform distribution or some heuristic approach. Then the solver $\pi$ performs an iterative, gradient-based optimization until a local optimum $x^{*}$ is found. The popular methods to conduct such gradient-based optimization in non-convex problems include the Interior Point Method (IPM) (Wright, 1997) and the Sequential Quadratic Programming (SQP) (Boggs & Tolle, 1995).
@@
As expressive generative models, the diffusion probabilistic models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020b) generate samples from random noises through an iterative denoising process. The forward process (i.e., diffusion process) iteratively adds Gaussian noise to sample $x_{k}$ at the $k$-th step according to the following equation:
@@
where $\widebar{\alpha}_{k}=\prod_{k^{\prime}=1}^{K}(1-\beta_{k^{\prime}})$ and $\varepsilon\sim\mathcal{N}(0,\mathbf{I})$ is a random sample from a standard Gaussian. The data generation process is the reverse denoising process that transforms random noise into data sample:
@@
with initial noisy data sampled from a random Gaussian distribution $x_{K}\sim\mathcal{N}(0,\mathbf{I})$. $p_{\theta}$ is the parameterized sampling distribution, and it can be optimized through an alternative $\varepsilon_{\theta}(x_{k},k)$ which is to predict $\varepsilon$ injected for $x_{k}$, for every diffusion step $k$. For conditional generation modeling, a conditional variable $y$ is added to both processes $q(x_{k+1}|x_{k},y)$ and $p_{\theta}(x_{k-1}|x_{k},y)$, respectively. The classifier-free guidance (Ho & Salimans, 2022) can be applied to further promote the conditional information, which learns both conditioned and unconditioned noise predictors as $\varepsilon_{\theta}(x_{k},k,y)$ and $\varepsilon_{\theta}(x_{k},k,\varnothing)$. Specifically, $\varepsilon_{\theta}$ is optimized with the following loss function:
@@ 4 Methodology
We first propose DiffuSolve, a general framework for efficient and robust non-convex trajectory optimization using diffusion models and NLP solvers. Then we introduce a novel constrained diffusion model specifically designed to guide the diffusion models to generate feasible solutions, which is included in DiffuSolve+.
4.1 DiffuSolve Framework
Figure 2: The DiffuSolve framework for solving non-convex trajectory optimization problems.
@@ 4.1 DiffuSolve Framework
It has been observed that for similar trajectory optimization problems $\mathcal{P}_{y}$ in Eq. (1), their solutions $x^{*}$ often exhibit similar structures (Amos et al., 2023; Li et al., 2023). Similar problems are those $\mathcal{P}_{y}$ with the same objective $J$ and constraints $g,h$, but with different parameters $y$. Thus, the key insight of DiffuSolve is to use a diffusion model to learn $p(x^{*}|y)$, the conditional probability distribution of locally optimal solutions based on pre-solved similar problems. Then the diffusion model can generalize and predict diverse solutions $\tilde{x}_{\text{new}}^{*}$ to new scenarios where the condition $y_{\text{new}}$ value is unseen. However, diffusion models inevitably make prediction errors, which can result in constraint violations. To address this, we use diffusion samples $\tilde{x}_{\text{new}}^{*}$ as initial guesses for the NLP solver, which fine-tunes them to derive the final solutions $x_{\text{new}}^{*}$, increasing both feasibility and optimality. This approach acts like a safety net, enhancing the robustness of the solutions.
@@
We illustrate our proposed DiffuSolve framework in Fig. 2. In the offline data generation process, we use an NLP solver $\pi$ to collect locally optimal and feasible solutions $x^{*}$ for various problem instances $\mathcal{P}_{y}$ with uniformly sampled initial guesses. The solutions with sub-optimal objective values are filtered out manually and the remaining dataset is used as the training data. For DiffuSolve training, we use a conditional diffusion model to learn $p(x^{*}|y)$ with classifier-free guidance (Ho & Salimans, 2022), with the loss function in Eq. 5. In the online testing process when a new problem $\mathcal{P}_{y_{\text{new}}}$ is presented with a priori unknown value $y_{\text{new}}$, we first use the diffusion model to predict diverse solution candidates $\tilde{x}_{\text{new}}^{*}$ as initial guesses. Then the NLP solver $\pi$ only requires minor adjustments to derive the final solutions $x_{\text{new}}^{*}$ with improved feasibility and optimality. This warm-starting approach is fully parallelizable for both diffusion sampling and the solving process of $\pi$.
4.2 DiffuSolve+ with Constrained Diffusion Model